# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v2

@@ Abstract
* Equal contribution. Work done primarily when the author was a postdoc researcher at Brown University.
@@ 1 Introduction
An autonomous agent operating in the real world must process low-level sensory and motor signals while reasoning about high-level objectives (Doncieux18; KONIDARIS20191). Analogous to how humans can perform complex tasks, like cooking or cleaning, without reasoning about muscle-level control, agents should have internal models of their skills that abstract away nuanced activities on the lower level. Such models must capture the necessary conditions for a skill to be executed (e.g., “pouring a teapot requires holding it first”) and the consequences of doing so (e.g., “pouring a teapot leaves it empty”). These two properties, known as preconditions and effects in the AI planning literature, enable compositional reasoning to identify long-horizon plans that can sequence lower-level skills to solve a task. Typically, these models must be specified manually. However, in real-world settings such skill representations may be nontrivial to acquire due to complex inter-skill constraints specific to the agent’s embodiment. This calls for algorithms that learn symbolic transition models of black-box skills without hand specification, enabling agents to directly utilize those skills to solve long-horizon tasks with off-the-shelf AI planners.
Traditional approaches of skill abstraction often require factorizing the low-level state space to learn classifiers for each symbolic representation, relying heavily on hand-collected transition data (Konidaris2018FromST). Recently, foundation models have enabled a new paradigm: generating semantically meaningful predicates directly from raw observations and directly evaluating their truth values on low-level observations (e.g., RGB images)—a process we refer to as generative predicate invention. Recent work has explored how foundation models can be used for predicate invention, by generating Python code to implement predicates (liang2025visualpredicator) or sampling large predicate pools followed by sub-selection. However, these methods produce ad-hoc planning representations that cannot be guaranteed to solve a given task, and leave core questions on predicate invention unanswered: what properties should these learned abstractions satisfy, and how can they be learned to achieve these properties?
@@ 2 Problem Setting
In this section, we briefly discuss our problem setting while defining it formally in Appendix A. We consider an agent equipped with a finite set of object-centric skills $\Omega$, modeled as black-box options (sutton1999options). The agent can execute any $\omega\in\Omega$ and determine whether it succeeds, but it does not possess a symbolic transition model of these skills. Without such a model, the agent cannot plan over a long horizon without reasoning at the low-level state space $\mathcal{S}$}, which is continuous, high-dimensional, and impractical for classical search. The goal of Skill Model Learning is to acquire a symbolic abstraction of skills that enables efficient composition via classical planning.
@@ 3.2 Predicate Invention
We now present our predicate invention algorithm, which, unlike prior work (silver2023predinvent; liang2025visualpredicator), does not require an initial set of predicates to bootstrap the invention process.
@@ 3.3 Operator Learning
Our operator learning procedure extends the associative model learning paradigm (arora2018review) to the setting of skill abstractions.
@@ 4.1 Implementation of Predicates
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 (openai_gpt5) for predicate generation and evaluation. In addition to the system performance reported in the main paper, we also conducted comprehensive studies of the component-wise reliability of the VLM in Appendix G.
@@ 4.2 Simulation
Figure 3: Screenshot of the Robotouille environment.
We first conduct experiments in Robotouille (gonzalez-pumariega2025robotouille), which is a simulated grid world kitchen domain with an agent that has five high-level skills: Pick, Place, Cut, Cook, and Stack. In the environment, there are several objects: a patty, lettuce, a top bun, and a bottom bun; there is also a cutting board and a stove for cutting the lettuce and cooking the patty, respectively. We design and categorize 50 abstract planning problems: 20 easy problems, whose optimal solutions have no more than 7 steps; 20 hard problems, whose optimal solutions have no more than 15 steps; and 10 impossible tasks that cannot be realized in the environment.
@@
  • ViLa (hu2023lookleapunveilingpower): This baseline is a closed-loop VLM-based approach that iteratively prompts a foundation model for the next action until the goal state is reached, given an image observation and the agent’s action history.
@@
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence consisting of 15 steps as their interaction budgets. We then evaluate each method on the evaluation set and report the average results in Table 1, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from liang2025visualpredicator). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, it is considered a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness. As shown in the table, SkillWrapper outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the system predicates baseline (Sys Preds.) on hard problems. Here we present the key insights, while leaving case studies, example operators, and failure modes and analysis of SkillWrapper in Appendix F.
@@ 4.3 Real Robots
To demonstrate the applicability of SkillWrapper for real-world agentic settings, we designed two sets of experiments with two robotic platforms: a Franka Emika Panda robot (Figures 4(a) and 4(b)) and a bimanual platform with two Kuka iiwa robots (Figures 4(c) and 4(d)). For both robot experiments, we assume that all skills are deterministic, which is a common assumption made by existing work (silver2023predinvent; han2024interpret; liang2025visualpredicator; athalye2025predicate). We give SkillWrapper 15 steps as its interaction budget per iteration. The supplementary material contains video demonstrations of both experiments.
@@ 5 Related Works
Our work uses pre-trained foundation models for learning symbolic representations of black-box robot skills useful to planning and close to human language and understanding. This work draws ideas from different fields of research such as model learning, abstraction learning, and task and motion planning (TAMP). Several methods have used foundation models (mainly LLMs) as high-level planners (saycan_corl; krishan2023sayplan; driess2023palme). Several approaches have used foundation models as robot action models (rt22023arxiv; shridhar23peract) or to generate reward functions for robot tasks (wang2024rlvlmf). Concurrent work has also explored how representations can be learned directly from pixels (athalye2025predicate). Although these approaches show promising results for short-horizon single-skill problems, they fail to scale to complex long-horizon problems (kambhampati2024llmscantplanhelp). Lastly, multiple approaches (han2024interpret; liang2025visualpredicator) have used foundation models to learn symbolic representations of robot skills, but require extensive feedback or prior knowledge from human experts. To the best of our knowledge, our work is the first to use a foundation model to automatically learn the human-interpretable symbolic characterization of robot skills with theoretical guarantees.
TAMP has long been used to solve complex robot tasks (dantam2018incremental; shah2020anytime; garrett2021integrated). However, these approaches require symbolic models of the robot skills for task planning. Various approaches have been developed to learn such symbolic models compatible with TAMP solvers from high-dimensional inputs (Konidaris2018FromST; silver2023predinvent; shah2024reals). Additionally, the abstract representations learned through these methods are not human-interpretable. On the other hand, we explicitly design our approach to work with high-dimensional inputs and generate human-interpretable abstractions using pre-trained foundation models. We consider abstractions as human-interpretable if they are semantically meaningful and use informative language descriptions. SkillWrapper also connects to other domains in robotics, and the full related work can be found in Appendix I.
@@ 7 Reproducibility Statement
To ensure reproducibility of our work, we provide source code and the prompts used in our experiments as supplementary materials. Although the reproducibility of real-world robot experiments is limited by hardware, simulation experiments run in Robotouille should be reliably reproduced, granted that the checkpoints of the foundation model (i.e., OpenAI’s GPT-5 (openai_gpt5)) have not been moved.
@@ A.1 Preliminaries
We model skills as object-centric options (sutton1999options) with discrete, object-typed parameters. Formally, a skill $\omega\in\Omega$ is defined by a tuple $({\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega},\Theta_{\omega}})$, where the initiation set $\mathcal{I}_{\omega}\subseteq\mathcal{S}$ contains states from which the skill may be executed; the policy $\pi_{\omega}$ controls the agent during the skill; the termination set $\beta_{\omega}\subseteq\mathcal{S}$ is the set of states at which the skill immediately terminates; and the skill parameters $\Theta_{\omega}=({\theta_{\omega}^{1},\dots,\theta_{\omega}^{k}})$ specify type constraints $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\mathcal{T}$ for valid skill arguments. Specifically, a skill $\omega\in\Omega$ may be instantiated to create a skill instance $\underline{\omega}$ using objects ${o}_{1:{k}}\in\mathcal{O}^{k}$ if and only if for $1\leq i\leq k$, $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\texttt{t}_{o_{i}}$. We assume that executing a skill only affects the state of objects passed as arguments. [^1] We do not, however, assume that a skill necessarily affects the state of all objects passed in as arguments.
@@
Because the agent must evaluate $\mathcal{I}_{\omega}$ on individual states, it cannot distinguish essential skill information from irrelevant details, making long-horizon planning combinatorially difficult. However, each skill affects only a few objects at once, leaving the rest of the world unchanged. The agent can exploit this property by instead learning a factored state representation—formalized here as first-order logic models in PDDL (mcdermott1998pddl)—providing an abstract transition model of its skills. Such abstractions enable the agent to use classical planning to compose its skills and accomplish unseen, long-horizon goals.
@@ B.1 Skill Sequence Proposal
Coverage ($C$).s
Coverage (Algorithm 3) evaluates the information gain on all possible pairs of consecutively executed skills over existing transitions after executing a new skill sequence. Specifically, the information gain is measured by the increase in Shannon entropy (shannon_entropy) over the distribution of all consecutive skill pairs resulting from executing the proposed skill sequence
@@ F.2 Result Analysis of Robotouille
We here present case studies of the failure modes of SkillWrapper and the baselines. One common failure mode is misclassifications during inference time, where we abstract low-level states specified in raw images into symbolic states using the predicates in the learned model. Since our contributions are more on the theory and algorithm design, we did not explore various potential techniques, such as ViperGPT (surís2023vipergptvisualinferencepython), that can possibly mitigate this issue.
@@ G.3 Other VLMs
We further examine the possibility of using open-source VLMs as alternatives for SkillWrapper. We choose Qwen3-VL-235B (Qwen3-VL) for comparison. To evaluate its capability, we conduct two sets of preliminary experiments: predicate classification and predicate invention.
@@ H.1 Planner and Planning Time
We use K* planner (katz-lee-socs2023) to generate top $K$ optimal plans, where $K$ in practice is the maximum planning budget. We use an i9-13900F CPU for running all the planning tasks. On average, each planning problem takes $0.0599$ seconds. Specifically, in Robotouille experiments, easy problems take $0.0549$ seconds, hard problems takes $0.0583$ seconds, and impossible problems take $0.0565$ seconds per problem; in Franka experiments, in-domain problems take $0.0529$ seconds, generalization problems take $0.5175$ seconds, and impossible tasks take $0.0516$ seconds per problem; in Bimanual Kuka experiments, all problems take $0.0553$ seconds on average.
@@ H.4 Robot Experiments
We employ a Franka Emika Research 3 robotic arm equipped with a UMI gripper (UMI). The workspace is observed by a single Intel RealSense D455 exocentric RGB-D camera, oriented to capture both the tabletop scene and the robot. The RGB data from this camera are used for learning symbolic models, while the depth information supports object pose estimation. Object poses are estimated using FoundationPose (wen2024foundationpose), which leverages high-fidelity 3D scanned models of the target objects. System-level communication and coordination are implemented in ROS 2 (Humble), which interfaces with motion planning, perception, and control modules. This setup supports five parameterized skills: Pick, Place, Stack, Pour, and Wipe. The first four skills (Pick, Place, Stack, and Pour) are executed through motion planning with the MoveIt framework, conditioned on both the end-effector and object poses. The Wipe skill is implemented by replaying a teleoperated trajectory.
@@
We use a robot with two horizontally mounted KUKA LBR iiwa 7 R800 manipulators, one with a BarrettHand BH8-282 gripper, and the other with a Schunk Dextrous Hand 2.0 gripper. The robot collects RGB data used for learning symbolic models with a MultiSense S7 camera mounted on a Pan-Tilt unit, while using an Intel RealSense D455 camera for RGB-D data used in pose estimation (wen2024foundationpose) of the objects in the scene. We use ROS 1 and KUKA FRI to communicate with the robot and utilize the built-in joint impedance control with position target as the low-level controller. At the high level, we create collision models of all objects in the scene and use a task and motion planner to generate motion plans for each skill. The Pick skills (compatible with knife and peanut butter jar) are implemented using motion planning. The OpenJar, Scoop, and Spread skills are implemented using a combination of motion planning and pre-defined trajectory playback.

### v3

@@ 1 Introduction
An autonomous agent operating in the real world must process low-level sensory and motor signals while reasoning about high-level objectives (Doncieux et al., 2018; Konidaris, 2019). Analogous to how humans can perform complex tasks, like cooking or cleaning, without reasoning about muscle-level control, agents should have internal models of their skills that abstract away nuanced activities on the lower level. Such models must capture the necessary conditions for a skill to be executed (e.g., “pouring a teapot requires holding it first”) and the consequences of doing so (e.g., “pouring a teapot leaves it empty”). These two properties, known as preconditions and effects in the AI planning literature, enable compositional reasoning to identify long-horizon plans that can sequence lower-level skills to solve a task. Typically, these models must be specified manually. However, in real-world settings such skill representations may be nontrivial to acquire due to complex inter-skill constraints specific to the agent’s embodiment. This calls for algorithms that learn symbolic transition models of black-box skills without hand specification, enabling agents to directly utilize those skills to solve long-horizon tasks with off-the-shelf AI planners.
Traditional approaches of skill abstraction often require factorizing the low-level state space to learn classifiers for each symbolic representation, relying heavily on hand-collected transition data (Konidaris et al., 2018). Recently, foundation models have enabled a new paradigm: generating semantically meaningful predicates directly from raw observations and directly evaluating their truth values on low-level observations (e.g., RGB images)—a process we refer to as generative predicate invention. Recent work has explored how foundation models can be used for predicate invention, by generating Python code to implement predicates (Liang et al., 2025) or sampling large predicate pools followed by sub-selection. However, these methods produce ad-hoc planning representations that cannot be guaranteed to solve a given task, and leave core questions on predicate invention unanswered: what properties should these learned abstractions satisfy, and how can they be learned to achieve these properties?
@@ 2 Problem Setting
In this section, we briefly discuss our problem setting while defining it formally in Appendix A. We consider an agent equipped with a finite set of object-centric skills $\Omega$, modeled as black-box options (Sutton et al., 1999). The agent can execute any $\omega\in\Omega$ and determine whether it succeeds, but it does not possess a symbolic transition model of these skills. Without such a model, the agent cannot plan over a long horizon without reasoning at the low-level state space $\mathcal{S}$}, which is continuous, high-dimensional, and impractical for classical search. The goal of Skill Model Learning is to acquire a symbolic abstraction of skills that enables efficient composition via classical planning.
@@ 3.2 Predicate Invention
We now present our predicate invention algorithm, which, unlike prior work (Silver et al., 2023; Liang et al., 2025), does not require an initial set of predicates to bootstrap the invention process.
@@ 3.3 Operator Learning
Our operator learning procedure extends the associative model learning paradigm (Arora et al., 2018) to the setting of skill abstraction.
@@ 4.1 Implementation of Predicates
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 (OpenAI, 2025) for predicate generation and evaluation. In addition to the system performance reported in the main paper, we also conducted comprehensive studies of the component-wise reliability of the VLM in Appendix G.
@@ 4.2 Simulation
Figure 3: Robotouille
environment.
We first conduct experiments in Robotouille (Gonzalez-Pumariega et al., 2025), which is a simulated grid world kitchen domain with an agent that has five high-level skills: Pick, Place, Cut, Cook, and Stack. In the environment, there are several objects: a patty, lettuce, a top bun, and a bottom bun; there is also a cutting board and a stove for cutting the lettuce and cooking the patty, respectively. We design and categorize 50 abstract planning problems: 20 easy problems, whose optimal solutions have no more than 7 steps; 20 hard problems, whose optimal solutions have no more than 15 steps; and 10 impossible tasks that cannot be realized in the environment.
@@
  • ViLa (Hu et al., 2023): This baseline is a closed-loop VLM-based approach that iteratively prompts a foundation model for the next action until the goal state is reached, given an image observation and the agent’s action history.
@@
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence consisting of 15 steps as their interaction budgets. We then evaluate each method on the evaluation set and report the average results in Table 1, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from Liang et al. (2025)). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, it is considered a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness.
As shown in the table, SkillWrapper outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the system predicates baseline (Sys Preds.) on hard problems. Here we present the key insights, while leaving case studies, example operators, and failure modes and analysis of SkillWrapper in Appendix F.
@@ 4.3 Real Robots
To demonstrate the applicability of SkillWrapper for real-world agentic settings, we designed two sets of experiments with two robotic platforms: a Franka Emika Panda robot (Figures 4(a) and 4(b)) and a bimanual platform with two Kuka iiwa robots (Figures 4(c) and 4(d)). For both robot experiments, we assume that all skills are deterministic, which is a common assumption made by existing work (Silver et al., 2023; Han et al., 2024; Liang et al., 2025; Athalye et al., 2025). We give SkillWrapper 15 steps as its interaction budget per iteration. The supplementary material contains video demonstrations of both experiments.
@@ 5 Related Works
Our work uses pre-trained foundation models for learning symbolic representations of black-box robot skills useful to planning and close to human language and understanding. This work draws ideas from different fields of research such as model learning, abstraction learning, and task and motion planning (TAMP). Several methods have used foundation models (mainly LLMs) as high-level planners (Ahn et al., 2022; Rana et al., 2023; Driess et al., 2023). Several approaches have used foundation models as robot action models (Brohan et al., 2023; Shridhar et al., 2023) or to generate reward functions for robot tasks (Wang et al., 2024b). Concurrent work has also explored how representations can be learned directly from pixels (Athalye et al., 2025). Although these approaches show promising results for short-horizon single-skill problems, they fail to scale to complex long-horizon problems (Kambhampati et al., 2024). Lastly, multiple approaches (Han et al., 2024; Liang et al., 2025) have used foundation models to learn symbolic representations of robot skills, but require extensive feedback or prior knowledge from human experts. To the best of our knowledge, our work is the first to use a foundation model to automatically learn the human-interpretable symbolic characterization of robot skills with theoretical guarantees.
TAMP has long been used to solve complex robot tasks (Dantam et al., 2018; Shah et al., 2020; Garrett et al., 2021). However, these approaches require symbolic models of the robot skills for task planning. Various approaches have been developed to learn such symbolic models compatible with TAMP solvers from high-dimensional inputs (Konidaris et al., 2018; Silver et al., 2023; Shah et al., 2024). Additionally, the abstract representations learned through these methods are not human-interpretable. On the other hand, we explicitly design our approach to work with high-dimensional inputs and generate human-interpretable abstractions using pre-trained foundation models. We consider abstractions as human-interpretable if they are semantically meaningful and use informative language descriptions. SkillWrapper also connects to other domains in robotics, and the full related work can be found in Appendix I.
@@ 7 Reproducibility Statement
To ensure reproducibility of our work, we provide source code and the prompts used in our experiments as supplementary materials. Although the reproducibility of real-world robot experiments is limited by hardware, simulation experiments run in Robotouille should be reliably reproduced, granted that the checkpoints of the foundation model (i.e., OpenAI’s GPT-5 (OpenAI, 2025)) have not been moved.
@@ A.1 Preliminaries
We model skills as object-centric options (Sutton et al., 1999) with discrete, object-typed parameters. Formally, a skill $\omega\in\Omega$ is defined by a tuple $({\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega},\Theta_{\omega}})$, where the initiation set $\mathcal{I}_{\omega}\subseteq\mathcal{S}$ contains states from which the skill may be executed; the policy $\pi_{\omega}$ controls the agent during the skill; the termination set $\beta_{\omega}\subseteq\mathcal{S}$ is the set of states at which the skill immediately terminates; and the skill parameters $\Theta_{\omega}=({\theta_{\omega}^{1},\dots,\theta_{\omega}^{k}})$ specify type constraints $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\mathcal{T}$ for valid skill arguments. Specifically, a skill $\omega\in\Omega$ may be instantiated to create a skill instance $\underline{\omega}$ using objects ${o}_{1:{k}}\in\mathcal{O}^{k}$ if and only if for $1\leq i\leq k$, $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\texttt{t}_{o_{i}}$. We assume that executing a skill only affects the state of objects passed as arguments. [^1] We do not, however, assume that a skill necessarily affects the state of all objects passed in as arguments.
@@
Because the agent must evaluate $\mathcal{I}_{\omega}$ on individual states, it cannot distinguish essential skill information from irrelevant details, making long-horizon planning combinatorially difficult. However, each skill affects only a few objects at once, leaving the rest of the world unchanged. The agent can exploit this property by instead learning a factored state representation—formalized here as first-order logic models in PDDL (McDermott et al., 1998)—providing an abstract transition model of its skills. Such abstractions enable the agent to use classical planning to compose its skills and accomplish unseen, long-horizon goals.
@@ B.1 Skill Sequence Proposal
Coverage ($C$).
Coverage (Algorithm 3) evaluates the information gain on all possible pairs of consecutively executed skills over existing transitions after executing a new skill sequence. Specifically, the information gain is measured by the increase in Shannon entropy (Shannon, 1948) over the distribution of all consecutive skill pairs resulting from executing the proposed skill sequence
@@ F.2 Result Analysis of Robotouille
We here present case studies of the failure modes of SkillWrapper and the baselines. One common failure mode is misclassifications during inference time, where we abstract low-level states specified in raw images into symbolic states using the predicates in the learned model. Since our contributions are more on the theory and algorithm design, we did not explore various potential techniques, such as ViperGPT (Surís et al., 2023), that can possibly mitigate this issue.
@@ G.3 Other VLMs
We further examine the possibility of using open-source VLMs as alternatives for SkillWrapper. We choose Qwen3-VL-235B (Bai et al., 2025) for comparison. To evaluate its capability, we conduct two sets of preliminary experiments: predicate classification and predicate invention.
@@ H.1 Planner and Planning Time
We use K* planner (Katz & Lee, 2023) to generate top $K$ optimal plans, where $K$ in practice is the maximum planning budget. We use an i9-13900F CPU for running all the planning tasks. On average, each planning problem takes $0.0599$ seconds. Specifically, in Robotouille experiments, easy problems take $0.0549$ seconds, hard problems takes $0.0583$ seconds, and impossible problems take $0.0565$ seconds per problem; in Franka experiments, in-domain problems take $0.0529$ seconds, generalization problems take $0.5175$ seconds, and impossible tasks take $0.0516$ seconds per problem; in Bimanual Kuka experiments, all problems take $0.0553$ seconds on average.
@@ H.4 Robot Experiments
We employ a Franka Emika Research 3 robotic arm equipped with a UMI gripper (Chi et al., 2024). The workspace is observed by a single Intel RealSense D455 exocentric RGB-D camera, oriented to capture both the tabletop scene and the robot. The RGB data from this camera are used for learning symbolic models, while the depth information supports object pose estimation. Object poses are estimated using FoundationPose (Wen et al., 2024), which leverages high-fidelity 3D scanned models of the target objects. System-level communication and coordination are implemented in ROS 2 (Humble), which interfaces with motion planning, perception, and control modules. This setup supports five parameterized skills: Pick, Place, Stack, Pour, and Wipe. The first four skills (Pick, Place, Stack, and Pour) are executed through motion planning with the MoveIt framework, conditioned on both the end-effector and object poses. The Wipe skill is implemented by replaying a teleoperated trajectory.
@@
We use a robot with two horizontally mounted KUKA LBR iiwa 7 R800 manipulators, one with a BarrettHand BH8-282 gripper, and the other with a Schunk Dextrous Hand 2.0 gripper. The robot collects RGB data used for learning symbolic models with a MultiSense S7 camera mounted on a Pan-Tilt unit, while using an Intel RealSense D455 camera for RGB-D data used in pose estimation (Wen et al., 2024) of the objects in the scene. We use ROS 1 and KUKA FRI to communicate with the robot and utilize the built-in joint impedance control with position target as the low-level controller. At the high level, we create collision models of all objects in the scene and use a task and motion planner to generate motion plans for each skill. The Pick skills (compatible with knife and peanut butter jar) are implemented using motion planning. The OpenJar, Scoop, and Spread skills are implemented using a combination of motion planning and pre-defined trajectory playback.